OpenStudy (nwolf):

As you prepare to move in to your new home, you are trying to decide whether to install a satellite dish TV system or purchase a DVD Player and rent DVDs. Here are the costs you were quoted: Satellite TV: One time installation charge of $48 and $38 per month. DVD: Purchase a DVD player for $160 and rent 8 DVDs per month at $3.00 per DVD. You want to compare the costs of each option over several months in order to make an appropriate decision. In order to decide which plan is a better deal, you do the following calculations: At 6 months the Satellite is a better deal because t

OpenStudy (anonymous):

satellite is always the best!

OpenStudy (nwolf):

I agree

OpenStudy (nwolf):

At 6 months the Satellite is a better deal because the plans will have cost you… Satellite: 48 + 38(6) = $276 DVD: 160 + 6(8)(3.00) = $304 At 12 months the DVD is a better deal because the plans will have cost you… Satellite: 48 + 38(12) = $504 DVD: 160 + 12(8)(3.00) = $448